# A -ray-emitting blazar B3 1239+376 at = 3.82 identified in a multi-wavelength context

### v1

@@ Abstract
Among thousands of extragalactic -ray emitters, only a handful of distant ( 3) sources are detected, yet they are cruial probes shedding light on the cosmic evolution of jets of active galactic nuclei and the initial phase of mass growth of supermassive black holes. Here, we report on a multi-band study of a radio quasar B3 1239+376 with = 3.82. By analyzing the Fermi-LAT data, a significant (globally 7.7) -ray source in its direction, with an estimated association probability of 0.91, is observed in a half-year period of 2025. The analyses also reveal the emergence of co-spatial -ray residues in prior epochs. Moreover, the -ray and infrared light curves are likely correlated, particularly, the two emissions climb to the peaking values at the same time. The temporal coincidence establishes a firm association relationship between the -ray source and the quasar. Therefore, B3 1239+376 is proposed as the third most distant -ray-emitting blazar to date. Benefiting from the multi-wavelength observations, broadband spectral energy distributions in different flux states are drawn and reproduced by the classic one-zone leptonic radiation model to investigate the jet properties. Considering the recent brightening in rays, prompt follow-up observations are encouraged, especially radio interferometry observations which may catch the potential ejection of a new jet blob.
@@ I Introduction
Active galactic nuclei (AGNs), are one of the most luminous sources of electromagnetic emissions in the universe, powered by the central supermassive black holes (SMBHs, Antonucci 1993; Urry & Padovani 1995). In the past 10 years, more than 400 quasars with z 5.7 have been discovered (Fan etĀ al., 2023). The existence of these objects suggests a prompt and eļ¬cient mass growth for the SMBHs, challenging the scenarios of SMBH formation in the early universe (Inayoshi etĀ al., 2020). The situation becomes more interesting when the high- AGNs with strong jets (i.e., jetted AGN) are detected (e.g., An & Romani, 2018; Belladitta etĀ al., 2020; Sbarrato etĀ al., 2022; BaƱados etĀ al., 2025). In this case, since the presence of relativistic jets is believed to relate with highly spinning accreting black holes (e.g., Tchekhovskoy etĀ al., 2011) that a high value for the radiation eļ¬ciency is expected (, Thorne 1974), a longer growth time of these sources is needed than that of the non-jetted AGNs.
@@
Blazars, including flat-spectrum radio quasars (FSRQs) and BL Lacertae objects (BL Lacs), are a subclass of AGNs that their strong jets are aligned with the line of sight and hence the beamed non-thermal jet emissions are overwhelming (Blandford etĀ al., 2019). A distinguished feature of the blazar radiation is the violent variability across diverse timescales (e.g., Ulrich etĀ al., 1997). The broadband jet emissions span from radio to very-high-energy -ray domains, characterized by a universal two-bump spectral shape in logālog plot. One bump is widely accepted from synchrotron emission, while the other one peaking in rays, is usually explained as inverse Compton (IC) by a population of relativistic electrons, which are also responsible for the synchrotron emission, scattering of soft photons from either inside (the synchrotron self-Compton, or SSC, Maraschi etĀ al. 1992) and/or outside (external Compton, or EC, Dermer & Schlickeiser 1993; Sikora etĀ al. 1994; BÅażejowski etĀ al. 2000) the jet. Alternatively, hadronic radiation models are adopted to describe the spatial and temporal coincidences between the incoming neutrino and the flaring blazars (e.g., IceCube Collaboration etĀ al., 2018; KM3NeT Collaboration etĀ al., 2025).
@@
In the -ray sky, the majority of the extragalactic sources are blazars (Ballet et al., 2023). Although the most distant blazar detected so far is with redshift as high as = 7 (Bañados et al., 2025), the redshift record of -ray-emitting blazars is 4.72 right now (i.e., B3 1428+422, Liao et al. 2018; Kreter et al. 2020; Gokus et al. 2025). Actually, there are only two blazars with -ray detections beyond redshift of 4, the other one is GB 1508+5714 ( = 4.3) (Ackermann et al., 2017; Liao et al., 2020; Gokus et al., 2024). The -ray radiations bring key information of the jet properties (Madejski & Sikora, 2016). In particular, -ray photons from high- blazars, are a powerful tool for probing the extragalactic background light (EBL, Fermi-LAT Collaboration et al. 2018). Since these sources are spectrally soft (Ackermann et al., 2017; Li et al., 2018) and the angular resolution of Fermi-LAT for the sub-GeV photons is worse than that of GeV photons, the major obstacle to detecting -ray signals from high- blazars is the limited localization. Catching correlated multi-wavelength activities between the -ray emission and those in other windows of electromagnetic radiation acts as a decisive evidence to determine the association relationship (e.g., Liao et al., 2019).
@@
B3 1239+376 is first known as a radio source observed by the Bologna Northern Cross Telescope (Colla etĀ al., 1973). It possesses a flat radio spectrum between 1.4 and 5 GHz (Taylor etĀ al., 1996) and is recognized as a high- quasar ( = 3.82, Vermeulen etĀ al. 1996), which makes it one of the brightest radio quasars in the early universe (i.e., 0.5 Jy). Although B3 1239+376 is not listed in the Roma-BZCat (hereafter BZCAT, Massaro etĀ al. 2009), the super-luminous motion of 5 has been detected (Britzen etĀ al., 2008). Meanwhile, its radio-loudness (RL=, Kellermann etĀ al. 1989) is calculated as high as 25,000, suggesting a blazar nature. In the X-ray perspective, a hard X-ray spectrum (i.e., = 1.5) is suggested, further supporting it as a blazar (McKeough etĀ al., 2016). In this study, multiwavelength data of B3 1239+376 are collected and analyzed, and its broadband temporal properties are reported (Section II), along with some discussions (Section III). Here we take a CDM cosmology with , , and (Planck Collaboration etĀ al., 2014).
@@ II.1 Fermi-LAT Data
We collected Fermi-LAT Pass 8 SOURCE data, with an energy range between 100 MeV and 500 GeV. The data were taken during the period from 2008 August 4 to 2025 August 4 (i.e., MJD 54683 - 60892), at a time length of 17-year. During the analysis, Fermitools software version 2.2.0 as well as Fermitools-data version 0.18 were adopted. A zenith angle cut (i.e., 90ā) along with the recommended quality-filter cuts (i.e., DATA_QUAL==1 && LAT_CONFIG==1) were included in the data filtering procedures. In the Unbinned likelihood analyses, in which -ray flux and spectrum were extracted by the gtlike task, the test statistic (TS = -2ln (/), Mattox etĀ al. 1996) was used to estimate the the significance of the -ray sources. In the definition, and denote the maximum likelihood values for the models with and without the target -ray source, respectively. Since the target is not embraced in an incremental version (data release 4) of the fourth Fermi-LAT catalog (4FGL-DR4, Ballet etĀ al. 2023), a test -ray sources at the radio position of B3 1239+376 was added in the source model. During the likelihood analysis, a region of interest (ROI) of 10ā centered on the target was considered. Parameters of the 4FGL-DR4 sources within the ROI, and those for the diffuse emission templates were set free, while others were frozen as the default values. In the temporal analysis, if detections of the background sources are marginal (i.e., TS 10), they were removed from the source model. In this case, for the target, 95% confidence level (CL) upper limits were calculated by the pyLikelihood UpperLimits tool, to replace the flux estimations.
@@
Since 4FGL-DR4 is based on the first 14-yr Fermi-LAT survey, an analysis for the additional 3-year data were performed. Interestingly, an excess in this direction emerged, with a TS value of 47 (locally 5.9). Therefore, a two-month time bin -ray light curve, was extracted to seek any potential activities recently. As shown in Figure 1, the background emission was dominant until the recent half-year period (i.e., MJD 60710 - 60892). Subsequent analysis aiming this specific half-year data revealed a significant residue (TS = 80, locally 8.2) towards B3 1239+376, see Figure 2. Here there are 34 trials (i.e., a selected half-year period for the entire 17-year dataset) and hence a global significance, after correction of the trial factor, was estimated as 7.7, suggesting that the detection is statistically significant. Assuming a simple power-law spectral template (i.e. , where is the photon index), the analysis yields a photon flux of ph . The source is spectrally soft, with an estimated photon index of . By comparison, the fit of the prior 2.5-year data only provided an upper limit (TS = 8), ph . Therefore, a strong variation with a 5-fold flux increase is indicated.
@@
Considering the limited angular resolution of Fermi-LAT, the possible influences from the background sources were checked. The closest background source is 4FGL J1243.2+3627, about 0.9ā away, however, it possesses a rather different spectral behavior (, Ballet etĀ al. 2023) than our target. We noted that the brightest background source in the ROI (i.e., 4FGL J1310.5+3221), which is 7.7ā from the ROI center, underwent a strong outburst in year 2025111https://fermi.gsfc.nasa.gov/ssc/data/access/lat/LightCurveRepository/source.html?source_name=4FGL_J1310.5+3221. Because the 95% CL containment angle of Fermi-LAT for 500 MeV photon is 222https://www.slac.stanford.edu/exp/glast/groups/canda/lat_Performance.htm, by selecting events with energies beyond 500Ā MeV, a further analysis was carried out. In this case, the contamination from the bright source could be significantly diminished. The detection remains to be significant, with a TS value of 39 which corresponds to a global significance of 4.7. Therefore, the new -ray source is likely intrinsic than artificially caused by the background sources. The obtained photon flux and index from the 500Ā MeV data analysis are consistent with that from the full data energy range.
@@
Localization analysis of the -ray source was followed. The optimized location was constrained as R.A. 190ā.6411 and decl. 37ā.2587, with a 95% CL uncertainty radius of 0.22ā. Corresponding angular separation between that and the radio position of B3 1239+376 is 0.11ā, and hence it falls into the -ray localization area, see Figure 2. When low energy threshold of the photons was set to 500Ā MeV, the best location is R.A. 190ā.4602 and decl. 37ā.2475. At this time, the angular separation (0.11ā) is still smaller than the a 95% CL uncertainty radius (0.21ā). Other potential low-energy association candidates were searched. Since blazars are the dominant population of the extragalactic sky, the blazar candidate catalogs were firstly looked into. No co-spatial sources listed in the BZCAT list (Massaro etĀ al., 2009) were found. Moreover, the result holds when the WIBRaLS2 catalog (DāAbrusco etĀ al., 2019) was investigated. Additionally, low-energy counterparts for -ray source have been searched in the high-frequency radio survey as well as from the VLBI observations (Abdollahi etĀ al., 2020). B3 1239+376 is the only one embraced by the CRATES catalog (Healey etĀ al., 2007) as well as the radio fundamental catalog (Petrov & Kovalev, 2025) in this direction. Adopting the Bayesian association method together with the corresponding prior probability value for CRATES catalog (0.33, Abdo etĀ al. 2010), the association probability was calculated as 0.91, above the threshold value of 0.8. Hence, B3 1239+376 is suggested to associate with the -ray source. Nevertheless, considering the relatively large localization uncertainty in the rays, detections of multiwavelength correlated emissions are needed to establish a firm association relationship.
@@
To provide a global picture of the temporal behavior of the -ray source, a 1-year time bin light curve for the entire 17-year dataset was built, as shown in Figure 3. Adopting the āvariability indexā test (Nolan etĀ al., 2012), the variability is suggested to be significant (). The recent brightening (i.e., the last time bin) stands out, and it has been picked up by the Bayesian block approach (Scargle etĀ al., 2013) indicative of a flare event, which is consistent with the results mentioned above. At the initial 9-year Fermi-LAT observation, the target was in a low flux state. The corresponding analysis yielded a marginal detection (TS =12), with a photon flux constrained to ph by setting the photon index as 2.6. It is noteworthy that there are tentative detections (TS 15) in two yearly time bins around year 2017 and 2020, respectively. Then a 4-year long quiescent state was followed, from which an upper limit of ph was obtained.
@@
Detailed investigations on the potential signals from year 2017 to 2020 were carried out. In the first step, a 1-month time bin light curve was generated, to identify the exact time range of the signals. Two epochs were selected, in which one is a 6-month period between MJD 58060 and 58242, and the other one lasting for 10 months is from MJD 58727 to 58969. The corresponding obtained TS values are 15 and 17, respectively. Since analyzing Fermi-LAT data in neither of the epochs alone yield significant detections (i.e., TS 25), a joint analysis of the two epochs was then performed. The joint analysis returned an increased TS value (TS = 31) and the enhanced statistics allowed us to localize the residue. The optimized location was gave as R.A. 190ā.642 and decl. 37ā.4649, along with a 95% CL uncertainty radius of 0.21ā. As shown in Figure 2, B3 1239+376 is included within the location error radius, since the angular separation is 0.15ā under this circumstance. The photon flux and the spectral information were also extracted. The former was obtained as ph , while the photon index was constrained as that is in agreement with the result from the flare in year 2025. Both the spatial and spectral information extracted in the joint analysis suggest the potential activities then likely also related to B3 1239+376.
II.2 Chandra X-ray Observatory data

### v2

# The -ray-emitting blazar B3 1239+376 at = 3.82 identified in a multi-wavelength context

@@ Abstract
Among thousands of extragalactic -ray emitters, only a handful of distant ( 3) sources are detected. Yet, they are crucial probes shedding light on the cosmic evolution of jets of active galactic nuclei and the initial phase of mass growth of supermassive black holes. Here, we report on a multi-band study of the radio quasar B3 1239+376 with = 3.82. By analyzing the Fermi-LAT data, a significant (globally 7.7) -ray source in its direction, with an estimated association probability of 0.91, is observed in a half-year period of 2025. The analysis also reveals the emergence of co-spatial -ray residues in prior epochs. Moreover, the -ray and infrared light curves obtained from WISE and SPHEREx observations are likely correlated, as we observe that the emissions in both bands peak at the same time. The temporal coincidence establishes a firm association relationship between the -ray source and the quasar. Therefore, B3 1239+376 is proposed as the third most distant -ray detected blazar to date. Benefiting from the multi-wavelength observations, broadband spectral energy distributions in different flux states are compiled and reproduced by the classic one-zone leptonic radiation model to investigate the jet properties. Considering the recent brightening in rays, prompt follow-up observations are encouraged, especially radio interferometry observations which may catch the potential ejection of a new jet blob.
@@ I Introduction
Active galactic nuclei (AGNs), are one of the most luminous sources of electromagnetic emission in the universe, powered by central supermassive black holes (SMBHs, Antonucci 1993; Urry and Padovani 1995). In the past 10 years, more than 400 quasars with z 5.7 have been discovered (Fan et al., 2023). The existence of these objects suggests a prompt and efficient mass growth for their SMBHs, challenging the scenarios of SMBH formation in the early universe (Inayoshi et al., 2020). The situation becomes more interesting when the high- AGNs with strong jets (i.e., jetted AGN) are detected (e.g., An and Romani, 2018; Belladitta et al., 2020; Sbarrato et al., 2022; Bañados et al., 2025). Jetted AGN usually harbor very heavy SMBHs, even in the early cosmic time (Ghisellini et al., 2010; Volonteri et al., 2011). The presence of relativistic jets is believed to relate with highly spinning accreting black holes (e.g., Tchekhovskoy et al., 2011), for which a high radiation efficiency is expected (, Thorne 1974). Therefore, a longer growth time for these SMBHs is needed compared for those residing in non-jetted AGNs, which indicates that the jets might play an important role in the initial phase of SMBH formation (Jolley and Kuncic, 2008).
@@
Blazars, which further distinguish into flat-spectrum radio quasars (FSRQs) and BL Lacertae objects (BL Lacs), are a subclass of AGNs for which their strong jets are aligned with our line of sight and hence the beamed non-thermal jet emission dominates the electromagnetic spectrum (Blandford et al., 2019). A distinguished feature of the blazar radiation is the violent variability across diverse timescales (e.g., Ulrich et al., 1997). The broadband jet emissions span from radio to very-high-energy -ray domains, characterized by a universal two-bump spectral shape in log–log plot. One bump is widely accepted from synchrotron emission, while the other one peaking in rays, is usually explained as inverse Compton (IC) by a population of relativistic electrons, which are also responsible for the synchrotron emission, scattering of soft photons from either inside (the synchrotron self-Compton, or SSC, Maraschi et al. 1992) and/or outside (external Compton, or EC, Dermer and Schlickeiser 1993; Sikora et al. 1994; Błażejowski et al. 2000) the jet. Alternatively, hadronic interactions, for instance the photo-hadronic () and hadronuclear () processes, are also raised to describe the high energy radiation of blazars (Mannheim, 1993; Mücke and Protheroe, 2001; Atoyan and Dermer, 2003). In addition to the electromagnetic radiation, decay of the pions generated from hadronic processes leads to the neutrino production as well. It is supported by recent multi-messenger observations that spatial and temporal coincidences between the incoming neutrino and the flaring blazars are occasionally detected (e.g., IceCube Collaboration et al., 2018; KM3NeT Collaboration et al., 2025).
@@
In the -ray sky, the majority of the extragalactic sources are blazars (Ballet et al., 2023). Although the most distant blazar detected so far is with redshift as high as = 7 (Bañados et al., 2025), the redshift record of -ray-emitting blazars is currently 4.72 (i.e., B3 1428+422, Liao et al. 2018; Kreter et al. 2020; Gokus et al. 2025). Actually, there are only two blazars with -ray detections beyond redshift of 4, the other one is GB 1508+5714 ( = 4.3) (Ackermann et al., 2017; Liao et al., 2020; Gokus et al., 2024). -ray radiation from AGN jets, exclusively produced by the non-thermal processes of the relativistic particles, possess a tight connection to the ultra-high-energy cosmic rays (Madejski and Sikora, 2016). Meanwhile, a large fraction of the electromagnetic radiation of blazars, especially the high- blazars, are released in the -ray domain (Ghisellini et al., 2011). Moreover, -ray photons from high- blazars, are a powerful tool for probing the extragalactic background light (EBL, Fermi-LAT Collaboration et al. 2018). Therefore, the rays bring key information on the AGN jets such as the acceleration and energy dissipation processes of emitting particles, and shed lights on the cosmic star formation history. However, since these -ray sources are spectrally soft (Ackermann et al., 2017; Li et al., 2018) and the angular resolution of Fermi-LAT for the sub-GeV photons is worse than that of GeV photons, the major obstacle to detecting -ray signals from high- blazars is the limited localization. Catching correlated multi-wavelength activities between the -ray emission and those in other windows of electromagnetic radiation acts as a decisive evidence to determine the association relationship (e.g., Liao et al., 2019).
@@
B3 1239+376 is first known as a radio source observed by the Bologna Northern Cross Telescope (Colla et al., 1973). It possesses a flat radio spectrum between 1.4 and 5 GHz (Taylor et al., 1996) and is recognized as a high- quasar ( = 3.82, Vermeulen et al. 1996), which makes it one of the brightest radio quasars in the early universe (i.e., 0.5 Jy). Although B3 1239+376 is not listed in the Roma-BZCat (hereafter BZCAT, Massaro et al. 2009), the super-luminous motion of 5 has been detected (Britzen et al., 2008). Meanwhile, its radio-loudness (RL=, Kellermann et al. 1989) is calculated as high as 25,000, suggesting a blazar nature. In the X-ray perspective, a hard X-ray spectrum (i.e., = 1.5) observed by Chandra X-ray observatory is suggested, further supporting it as a blazar (McKeough et al., 2016). In this study, multiwavelength data of B3 1239+376 are collected and analyzed, and its broadband temporal properties are reported (Section II), along with some discussions (Section IV). Here we consider a CDM cosmology with , , and (Planck Collaboration et al., 2014).
@@ II.1 Fermi-LAT Data
We collected Fermi-LAT Pass 8 SOURCE data, with an energy range between 100 MeV and 500 GeV. The data time range is from 2008 August 4 to 2025 August 4 (i.e., MJD 54683 - 60892), covering a total of 17 years. During the analysis, we adopted the Fermitools software version 2.2.0 as well as Fermitools-data version 0.18. A zenith angle cut (i.e., 90∘) along with the recommended quality-filter cuts (i.e., DATA_QUAL==1 && LAT_CONFIG==1) were included in the data filtering procedures. In the Unbinned likelihood analyses, in which -ray flux and spectrum were extracted by the gtlike task, we used the test statistic (TS = -2ln (/), Mattox et al. 1996) to estimate the the significance of the -ray sources. By definition, and denote the maximum likelihood values for the models with and without the target -ray source, respectively. Since the target is not included in any version of the Fermi-LAT catalogs, including the most recent incremental version (i.e., data release 4 of the fourth Fermi-LAT catalog, or 4FGL-DR4, Ballet et al. 2023), we added a test -ray source at the radio position of B3 1239+376 in the source model. We chosen power-law function to describe the spectral distribution of the test source, of which the initial spectral index was set as 2.5. During the likelihood analysis, all 4FGL-DR4 sources within 15∘ of the target were considered. Spectral and normalization parameters of the sources lying within 10∘ radius of interest (ROI), together with normalization parameters of the diffuse emission templates (i.e., gll_iem_v07 and iso_P8R3_SOURCE_V3_v1), were set free. On the other hand, for sources outside the ROI, we frozen their parameters as the default 4FGL-DR4 values. In the temporal analysis, if detections of the background sources are marginal (i.e., TS 10), we removed them from the source model. In this case, for the target, 95% confidence level (CL) upper limits were calculated by the pyLikelihood UpperLimits tool, to replace the fluxes. Here we took a linear binning approach, in other words, length of each time bin is equal. Since the variability timescales of blazars are diverse, across minutes to years, different binning strategies were adopted corresponding to different cases. The details were introduced in the following context.
@@
Since data from the first 14-yr Fermi-LAT survey (i.e., from 2008-08-04 to 2022-08-02) had been already analyzed and the results were released in 4FGL-DR4 (Ballet et al., 2023), at first, we focused on data from the recent observations (i.e., from 2022-08-02 to 2025-08-04). Interestingly, an excess in this direction emerged in our analyses, with a TS value of 47 (locally 5.9). Therefore, a two-month time bin -ray light curve, was extracted to seek any potential activities recently. As shown in Figure 1, no significant -ray signals were seen until the recent half-year period (i.e., MJD 60710 - 60892). Subsequent analysis aiming this specific half-year data revealed a significant residue (TS = 80, locally 8.2) towards B3 1239+376, see Figure 2. Here there are 34 trials (i.e., a selected half-year period for the entire 17-year dataset) and hence a global significance, after correction of the trial factor, was estimated as 7.7, suggesting that the detection is statistically significant. Assuming a simple power-law spectral template (i.e. , where is the photon index), the analysis yields a photon flux of ph . The source is spectrally soft, with an estimated photon index of . By comparison, the fit of the prior 2.5-year data only provided an upper limit (TS = 8), ph . Therefore, a strong variation with a 5-fold flux increase is indicated.
@@
Considering the limited angular resolution of Fermi-LAT, we checked the possible influences from the background sources. The closest background source is 4FGL J1243.2+3627, about 0.9∘ away, however, it possesses a rather different spectral behavior (, Ballet et al. 2023) than our target. We noted that the brightest background source in the ROI (i.e., 4FGL J1310.5+3221, 2.23, Ballet et al. 2023), which is 7.7∘ from the ROI center, underwent a strong outburst in year 2025111https://fermi.gsfc.nasa.gov/ssc/data/access/lat/LightCurveRepository/source.html?source_name=4FGL_J1310.5+3221. Because the 95% CL containment angle of Fermi-LAT for 500 MeV photon is 222https://www.slac.stanford.edu/exp/glast/groups/canda/lat_Performance.htm, by selecting events with energies beyond 500 MeV, we carried out a further analysis. In this case, the contamination from the bright source could be significantly diminished. The detection remains to be significant, with a TS value of 39 which corresponds to a global significance of 4.7. Therefore, the new -ray source is likely intrinsic than artificially caused by the background sources. The obtained photon flux and index from the 500 MeV data analysis are consistent with that from the full data energy range.
@@
Then we performed the localization analysis of the -ray source. The optimized location was constrained as R.A. 190∘.6411 and decl. 37∘.2587, with a 95% CL uncertainty radius of 0.22∘. Corresponding angular separation between that and the radio position of B3 1239+376 is 0.11∘, and hence it falls into the -ray localization area, see Figure 2. When low energy threshold of the photons was set to 500 MeV, the best location is R.A. 190∘.4602 and decl. 37∘.2475. At this time, the angular separation (0.11∘) is still smaller than the 95% CL uncertainty radius (0.21∘). We also searched other potential low-energy association candidates. Since blazars are the dominant population of the extragalactic sky, we firstly looked into the blazar candidate catalogs. No co-spatial sources were found in either the BZCAT list (Massaro et al., 2009) were found nor the WIBRaLS2 catalog (D’Abrusco et al., 2019) was investigated. Additionally, we investigated low-energy counterparts for -ray source by searching the high-frequency radio survey as well as from the VLBI observations (Abdollahi et al., 2020). No other sources than B3 1239+376 in the CRATES catalog were found (Healey et al., 2007), so as the radio fundamental catalog (Petrov and Kovalev, 2025) in this direction. Adopting the Bayesian association method together with the corresponding prior probability value for CRATES catalog (0.33, Abdo and others 2010), we calculated the association probability as 0.91, above the threshold value of 0.8. Hence, B3 1239+376 is suggested as the counterpart to the -ray source. Nevertheless, considering the relatively large localization uncertainty in the rays, detections of multiwavelength correlated emissions are needed to establish a firm association relationship.
@@
To provide a global picture of the temporal behavior of the -ray source, we built a 1-year time bin light curve for the entire 17-year dataset, as shown in Figure 3. Adopting the “variability index” test (Nolan et al., 2012), the variability is suggested to be significant (). The recent brightening (i.e., the last time bin) stands out, and it has been picked up by the Bayesian block approach (Scargle et al., 2013) indicative of a flare event, which is consistent with the results mentioned above. At the initial 9-year Fermi-LAT observation, the target was in a low flux state. The corresponding analysis yielded a marginal detection (TS =12), with a photon flux constrained to ph by setting the photon index as 2.6. It is noteworthy that there are tentative detections (TS 15) in two yearly time bins around year 2017 and 2020, respectively. Then a 4-year long quiescent state was followed, from which an upper limit of ph was obtained.
@@
We carried out detailed investigations on the potential signals from year 2017 to 2020. In the first step, we extracted a 1-month time bin light curve, to identify the exact time range of the signals. We selected two epochs, in which one is a 6-month period between MJD 58060 and 58242, and the other one lasting for 10 months is from MJD 58727 to 58969. The corresponding TS values are 15 and 17, respectively. Since analyzing Fermi-LAT data in neither of the epochs alone yield significant detections (i.e., TS 25), we performed a joint analysis of the two epochs then. The joint analysis returned an increased TS value (TS = 31) and the enhanced statistics allowed us to localize the residue. The optimized location was given as R.A. 190∘.642 and decl. 37∘.4649, along with a 95% CL uncertainty radius of 0.21∘. As shown in Figure 2, B3 1239+376 is included within the location error radius, since the angular separation is 0.15∘ under this circumstance. We also derived the photon flux and the spectral information. The former was obtained as ph , while the photon index was constrained as that is in agreement with the result from the flare in year 2025. Both the spatial and spectral information extracted in the joint analysis suggest the potential activities then likely also related to B3 1239+376.
II.2 X-ray data
II.2.1 Chandra data